OpenStudy (anonymous):

if arcsinA=5/8 , what does SinA =?

OpenStudy (jdoe0001):

\(\bf sin^{-1}(A)=\cfrac{5}{8}\implies sin\left[sin^{-1}( {\color{brown}{ A }}) \right]=sin\left( \cfrac{5}{8}\right) \\ \quad \\ recall\implies sin[sin^{-1}({\color{brown}{ 123}})]={\color{brown}{ 123}}\qquad sin^{-1}[sin^{}({\color{brown}{ \theta}})]={\color{brown}{ \theta}}\)

OpenStudy (jdoe0001):

well...hmmm you don't really need to take sin() to both sides anyhow though you needed the "A" value but notice that line below anyhow

OpenStudy (anonymous):

so the answer is 5/8?

OpenStudy (jdoe0001):

well... close not quite we first off, want to know ... something keep in mind that a inverse trig function, "takes a value" and "returns an angle" so

OpenStudy (jdoe0001):

\(\bf sin^{-1}(A)=\cfrac{5}{8}\implies sin\left[sin^{-1}( {\color{brown}{ A }}) \right]=sin\left( \cfrac{5}{8}\right) \\ \quad \\ recall\implies sin[sin^{-1}({\color{brown}{ 123}})]={\color{brown}{ 123}}\qquad sin^{-1}[sin^{}({\color{brown}{ \theta}})]={\color{brown}{ \theta}} \\ \quad \\ sin\left[sin^{-1}( {\color{brown}{ A }}) \right]=sin\left( \cfrac{5}{8}\right)\implies A=sin\left( \cfrac{5}{8}\right) \\ \quad \\ \textit{what's the }\ sin(A)?\qquad well \\ \quad \\ sin(A)=sin\left[ sin\left( \cfrac{5}{8}\right) \right]\)

OpenStudy (jdoe0001):

so the inverse function took the value returned an angle sine of that angle is, well, whatever was inside the inverse function that is value A so to find what the sine of THAT value A is we take out A, firstly then take its sine

OpenStudy (anonymous):

huh? its sina= sin(sin(5/8))???

OpenStudy (jdoe0001):

yes, because, notice above given value \(\Large \bf A\) is just a value

OpenStudy (jdoe0001):

hehe....see it.? no? the value A is inside the inverse function we have to take it out to know what it's once it's outside we can just take the sine of it